# mycap[Listing][List of Listings]

### v1

@@ Abstract
Audio often serves as an auxiliary modality in video understanding tasks of audio-visual large language models (LLMs), merely assisting in the comprehension of visual information. However, a thorough understanding of videos significantly depends on auditory information, as audio offers critical context, emotional cues, and semantic meaning that visual data alone often lacks. This paper proposes an audio-centric video understanding benchmark (ACVUBench) to evaluate the video comprehension capabilities of multimodal LLMs with a particular focus on auditory information. Specifically, ACVUBench incorporates 2,662 videos spanning 18 different domains with rich auditory information, together with over 13k high-quality human annotated or validated question-answer pairs. Moreover, ACVUBench introduces a suite of carefully designed audio-centric tasks, holistically testing the understanding of both audio content and audio-visual interactions in videos. A thorough evaluation across a diverse range of open-source and proprietary multimodal LLMs is performed, followed by the analyses of deficiencies in audio-visual LLMs. Demos are available at https://github.com/lark-png/ACVUBench.
@@ 1 Introduction
ACVUBench 2,662 67.8 13,774 8 M+SA
Table 1: Comparison of ACVUBench with other representative video and audio-visual benchmarks. “Annotation” is the annotation type (A: Automatic, SA: Semi-Automatic, M: Manual, M+SA: Mixed); “w. Audio” indicates the presence of audio; and “Ori. Videos” denotes whether the videos are newly collected () or repurposed ().
@@
However, a comprehensive understanding of videos relies heavily on auditory information, as audio, especially speech, provides essential context, emotional cues, and semantic meaning often absent from visual data alone. The lack of comprehensive assessment of audio processing capabilities significantly hinders the progress of MLLMs towards general video understanding. Consequently, there is an urgent need for dedicated efforts to evaluate model capabilities to understand and process speech and audio in videos. Moreover, existing audio-visual benchmarks often lack detailed human annotations specifically designed to assess audio-centric understanding. To address this gap, this paper proposes ACVUBench, the first comprehensive benchmark specifically designed for audio-centric video understanding. The key differences of ACVUBench compared with examples of existing benchmarks are shown in Table 1.
ACVUBench introduces a suite of carefully designed audio-centric video understanding tasks, which cover not only content understanding of the video but also the interaction between audio and visual information, such as time synchronization or content matching. Besides, ACVUBench adopts a dual-dataset design which combines high-quality human annotation with scalable, semi-automatic annotation to obtain a holistic evaluation benchmark. Specifically, a novel Gemini-powered in-context learning data creation approach is proposed to generate high-quality question-answer pairs that are further validated by human annotators. ACVUBench incorporates 2,662 videos meticulously selected from YouTube, covering 18 audio-centric domains. This collection is split into two partitions: (1) AV-Human, comprising 698 videos with 2,100 expertly human-annotated question-answer pairs; and (2) AV-Gemini, containing the remaining 1,964 videos with 11,674 question-answer pairs obtained from the proposed in-context learning data creation method.
A thorough evaluation of a diverse range of state-of-the-art MLLMs is performed on ACVUBench, encompassing audio-visual, visual-only, and audio-only LLMs. In addition to standard accuracy, ACVUBench also incorporates a confidence-thresholded accuracy metric based on the first-token probability to exclude random guesses that sometimes occur in multiple choice questions (MCQs), revealing the true capabilities of models. Analyses and insights into the strengths and weaknesses of different model types are provided across different audio-centric tasks. The main contributions are summarized as follows:
  • We introduce ACVUBench, the first comprehensive benchmark for Audio-Centric Video Understanding, featuring diverse tasks encompassing both audio content understanding and audio-visual alignment.
  • ACVUBench adopts a large-scale semi-automatic annotation process for 2,662 videos spanning 18 audio-centric domains by adopting an in-context learning-based data-creating approach powered by Gemini 1.5 Pro.
  • A thorough evaluation of a diverse range of audio-visual, visual-only, and audio-only models is conducted on ACVUBench, revealing their capabilities in understanding audio in videos.
@@ 2 Related Work
Early research in multimodal alignment explored the creation of shared representations across images, audio, and text Aytar et al. (2017), demonstrating the potential of transferable classifiers. Further research on shared representations across modalities leads to encoders that can align with LLMs like Whisper Radford et al. (2023) or CLIP Radford et al. (2021), making MLLMs technically possible. Recent progress in audio-visual learning has spurred the development of large language models (LLMs) capable of processing both audio and visual information. Early video understanding works like VideoChat Li et al. (2023a) uses external subtitles for speech. Subsequent models have explored approaches to fuse audio and visual modalities. PandaGPTSu et al. (2023) integrates ImageBindGirdhar et al. (2023) and VicunaChiang et al. (2023), leveraging image-text pairings for enhanced cross-modal understanding, while VASTChen et al. (2023b), trained on the large-scale VAST-27MChen et al. (2023b) dataset, further advances multimodal capabilities. Models like CATYe et al. (2025) employ specialized architectures, such as clue aggregators, to effectively process audio-visual events. Video-LLaMAZhang et al. (2023), VideoLLaMA2Cheng et al. (2024) and Baichuan OmniLi et al. (2024b) have been fine-tuned specifically on audio data to improve multimodal comprehension. Architectures like the multi-resolution causal Q-former in video-SALMONNSun et al. (2024b) aim to bridge pre-trained audio-visual encoders with LLMs efficiently. Furthermore, models like NExTGPTWu et al. (2024), OneLLMHan et al. (2024), and VITAFu et al. (2024b) use “omni-encoders” to uniformly process various data modalities, including audio and video.
@@
Numerous efforts have focused on developing video benchmarksXiao et al. (2021); Li et al. (2021, 2024a); Ning et al. (2023); Li et al. (2023b); Liu et al. (2024); Mangalam et al. (2023); Fu et al. (2024a); Fang et al. (2024), predominantly targeting the visual aspects of video content. However, the crucial role of audio in video understanding has received comparatively less attention. Several early datasets explore the use of audio in video understanding. AVSDAlamri et al. (2019) focuses on scene-aware dialogue, Pano-AVQAYun et al. (2021) on panoramic videos, MUSIC-AVQALi et al. (2022) specifically on musical contexts, and VGGSoundChen et al. (2020) on short videos of audio events. Thus, these datasets target specific scenarios. While AVQAYang et al. (2022) explores the synergy between audio and visual information but is limited by shorter video lengths and simplistic task design. More recent datasets like UnAV-100Geng et al. (2023), VALORChen et al. (2023a), and VAST-27MChen et al. (2023b) offer large-scale data but prioritize specific tasks, such as audio event detection, retrieval, and captioning. Similarly, LongVALEGeng et al. (2024) focuses on audio-visual-language events, and AVHBenchSung-Bin et al. (2024) on cross-modal hallucination. These benchmarks, often lacking human annotation, are less suited for evaluating the nuanced and complex aspects of audio-centric understanding, particularly alignment capabilities, targeted by ACVUBench.
@@ 3.1 Data Curation Principles
Focusing on Audio-Centric Understanding in Video. Unlike existing video benchmarks that predominantly target visual understanding, ACVUBench focuses specifically on the crucial role of audio in video comprehension. To this end, the tasks in ACVUBench are designed to focus on the following two aspects: (1) Audio Content Understanding, requiring models to perceive and understand audio information; (2) Audio Visual Alignment, requiring the models to align audio with the corresponding visual information.
Covering Diverse Audio-Centric Video Domains and Comprehensive Annotation. To thoroughly evaluate audio-centric understanding capabilities in MLLMs, ACVUBench incorporates a diverse selection of 18 video domains where audio plays a central role. For robust evaluation, ACVUBench offers both AV-Human, a meticulously human-annotated dataset, and AV-Gemini, a significantly larger dataset generated via a novel Gemini-augmented approach. This combination of diverse content and rich annotation enables ACVUBench to effectively support the research in audio-centric video understanding.
@@ 3.2 Audio-Centric Task Definition
Table 2: Task examples of ACVUBench. “Audio Visual Segment Matching” is to sort four randomly shuffled audio segments in a video into their original order. “Audio Visual Speaker Diarization” is to transcribe the speech of a visually specified speaker with the Word Error Rate (WER) as the metric.
@@ 3.3 Dataset Construction
ACVUBench contains 2,662 videos with 13,774 questions in total with each question either created or validated by human annotators. This section describes the data construction in detail.
@@ 3.3.1 Video Collection
Figure 1: The 18 video domains of ACVUBench from which the videos are collected.
We define 18 audio-centric video domains as shown in Figure 1 to cover a diverse range of scenarios in which audio plays a crucial role in comprehension. Our annotators searched YouTube for videos within these audio-centric domains to manually collect a total of 2,662 raw videos, as depicted in Figure 2. The collected videos were required to be in English, rich in audio content, and recently published (mostly after 2023) to minimize overlap with existing training sets. Videos were intentionally limited to two minutes as driven by the following considerations: First, the nature of our audio-centric tasks, which focus on analyzing specific audio events or aligning audio cues with corresponding visual information within localized segments, lends itself well to evaluation within this shorter time frame. Second, most popular open-source MLLMs struggle to process longer videos, and hence short videos allow us to perform a wider range of comparisons. Finally, extending the length of the video would lead to an exponential increase in the cost and potential inaccuracies of manual annotation, without significantly enhancing the discriminatory power of our benchmark. These videos further went through a validation process, where independent annotators confirmed the designated domains and the presence of sufficient and relevant audio information. The detailed video collection process is provided in Appendix D.
@@ 3.3.2 Human Annotation Partition
Figure 2: The Pipeline of ACVUBench. The pipeline consists of three main parts: video collection, human annotation and in-context learning creation. Annotators select videos from YouTube and these videos are then used for human annotation, where question-answer pairs are designed based on various task types. The remaining videos are used for in-context learning with Gemini 1.5 Pro. The in-context learning process involves four stages: (1) Demonstration Selection; (2) Prompt Designing; (3) Generation; (4) Refining.
@@
Annotation Procedure.
We selected 698 videos for full annotation where annotators were given detailed task definitions and annotation examples. Then, for each video, annotators selected the three most suitable task types and created corresponding multiple-choice questions and answer options, resulting in a total of 2,100 question-answer pairs. For AVDiar, annotators selected 260 videos from the full annotation set that contain a higher number of speakers (3.59 speakers per video on average). Detailed annotations for these videos included speaker appearance, start and end times of utterances, and speech transcriptions. The Audio-Visual Matching task did not require manual annotation, as audio-video segment pairings were generated automatically using a script.
Quality Assurance.
For each task, annotation experts evaluated and selected suitable annotators based on task-specific requirements (e.g., strong listening skills for speaker diarization). Selected annotators underwent online training and trial annotation overseen by dedicated managers. Upon demonstrating proficiency, annotators began formal annotation. A multi-stage quality control process ensured high-quality annotations. First, dedicated quality control specialists conducted an initial review and revision. Subsequently, a separate group of annotators performed a complete cross-validation check. Only data passing both checks was accepted. During annotation, special attention was paid to ensuring questions avoided common sense reasoning, required multimodal understanding (not answerable from text alone), and aligned precisely with the intended task type. This rigorous process ensured the accuracy, consistency, and overall quality of the ACVUBench annotations.
@@ 3.3.3 Gemini-Augmented Partition
Stage 3: Automated Annotation Generation. We provided Gemini with the remaining 1,964 raw videos from our collected set of 2,662. For each video and each task type, Gemini generates a question, four options and the correct option, resulting in a total of 11,674 generated QA pairs.
Stage 4: Human Review and Refinement. After generation, human annotators meticulously reviewed all QA pairs by focusing on three key aspects: (1) The accuracy and relevance of the generated questions and options; (2) The adherence to the specific requirements of each task type; and (3) The correctness of the generated answer. Any identified discrepancies, such as incorrect answers, were manually corrected by the annotators. Correction examples are shown in Appendix F.
@@ 4 Experiment
Models AIE (%) ACC (%) AEL (%) AVCM (%) AVOM (%) AVTM (%) Audio Visual MLLMs Gemini 1.5 Pro (2024)     90.02 / 94.70 50.85 / 39.99 83.43 / 73.32 74.05 / 83.13 74.79 / 76.62 81.22 / 83.68 VideoLLaMA2 (7B) (2024)     42.30 / 35.65 32.77 / 32.88 32.02 / 25.75 46.80 / 57.41 49.60 / 46.11 48.92 / 40.90 video-SALMONN (13B) (2024b)     52.45 / 32.86 33.33 / 30.17 36.05 / 32.48 34.47 / 41.42 30.25 / 31.42 39.95 / 36.21 PandaGPT (13B) (2023)     24.94 / 30.82 25.00 / 29.13 26.16 / 30.61 24.31 / 27.54 26.28 / 28.44 25.77 / 26.36 Visual MLLMs GPT-4o (2024)     60.37 / 61.43 35.00 / 31.14 33.14 / 34.57 51.81 / 55.71 59.34 / 55.43 70.69 / 61.71 Qwen2-VL (7B) (2024)     59.21 / 53.48 41.67 / 33.18 34.88 / 39.44 62.90 / 65.08 55.85 / 57.01 69.74 / 58.37 LLaVA-Video (7B) (2024)     54.19 / 55.77 39.05 / 30.62 24.83 / 38.17 62.83 / 62.53 61.87 / 58.21 62.94 / 60.04 InternVL2 (76B) (2024)     53.15 / 45.49 25.83 / 26.83 30.23 / 26.89 58.42 / 57.79 53.18 / 50.13 61.70 / 53.25 InternVL2 (8B) (2024)     44.52 / 37.70 30.00 / 25.24 29.07 / 26.58 49.25 / 53.54 47.43 / 43.38 53.19 / 43.69 VILA-1.5 (8B) (2024)     41.72 / 38.28 28.33 / 26.78 28.49 / 36.61 53.30 / 62.38 46.41 / 47.77 46.34 / 46.64 VideoLLaVA (7B) (2023)     34.73 / 27.97 15.00 / 19.10 22.67 / 17.30 36.67 / 42.36 34.09 / 35.86 35.93 / 29.69 Audio MLLM SALMONN (13B) (2024)     45.92 / 45.18 33.33 / 31.68 29.65 / 31.56 32.62 / 32.70 36.76 / 33.11 34.52 / 38.39
Table 4: Overall Performance on ACVUBench. This table presents the overall performance of each model on the AV-Human, AV-Gemini, and Overall (combined AV-Human and AV-Gemini) datasets. Performance is measured by accuracy (%) on MCQs.
Models AV-Human AV-Gemini Overall Audio Visual MLLMs Gemini 1.5 Pro (2024)     78.34     75.21     75.67 VideoLLaMA2 (7B) (2024)     44.90     39.75     40.56 video-SALMONN (13B) (2024b)     38.33     34.08     34.74 PandaGPT (13B) (2023)     25.38     28.83     28.31 Visual MLLMs GPT-4o (2024)     56.62     50.00111We evaluated GPT-4 on a subset of 2100 questions randomly sampled from the full question set with balanced task types, sufficiently representing the original task distribution.     53.31 Qwen2-VL (7B) (2024)     58.38     51.05     52.26 LLaVA-Video (7B) (2024)     56.52     50.62     51.48 InternVL2 (76B) (2024)     52.62     43.29     44.72 InternVL2 (8B) (2024)     45.90     38.31     39.47 VILA-1.5 (8B) (2024)     44.48     43.05     43.27 VideoLLaVA (7B) (2023)     33.14     28.70     29.37 Audio MLLM SALMONN (13B) (2024)     36.48     35.43     35.59

### v3

@@ Abstract
Audio often serves as an auxiliary modality in video understanding tasks of audio-visual large language models (LLMs), merely assisting in the comprehension of visual information. However, a thorough understanding of videos significantly depends on auditory information, as audio offers critical context, emotional cues, and semantic meaning that visual data alone often lacks. This paper proposes an audio-centric video understanding benchmark (AVUT) to evaluate the video comprehension capabilities of multimodal LLMs with a particular focus on auditory information. AVUT introduces a suite of carefully designed audio-centric tasks, holistically testing the understanding of both audio content and audio-visual interactions in videos. Moreover, this work points out the text shortcut problem that largely exists in other benchmarks where the correct answer can be found from question text alone without needing videos. AVUT addresses this problem by proposing a answer permutation-based filtering mechanism. A thorough evaluation across a diverse range of open-source and proprietary multimodal LLMs is performed, followed by the analyses of deficiencies in audio-visual LLMs. Demos and data are available at https://github.com/lark-png/AVUT.
@@ 1 Introduction
AVUT 2,662 67.8 11,609 8 M+SA
Table 1: Comparison of AVUT with other representative video and audio-visual benchmarks. “Annotation” is the annotation type (A: Automatic, SA: Semi-Automatic, M: Manual, M+SA: Mixed); “w. Audio” indicates the presence of audio; and “Ori. Videos” denotes whether the videos are newly collected () or repurposed ().
@@
However, a comprehensive understanding of videos relies heavily on auditory information, as it provides essential context, emotional cues, and semantic meaning often absent from visual data alone. Existing benchmarks either overlooks the importance of audio in video understanding, or only focus on the speech content and ignoring other information. Consequently, there is an urgent need for dedicated efforts to evaluate model capabilities to understand and process speech and audio in videos. Moreover, most existing video understanding benchmarks suffer from the text shortcut problem, where the model can deduce the correct answer by simply focusing on the text alone. To fill the gaps, we propose AVUT, the text-shortcut-free benchmark specifically designed for audio-centric video understanding. The key differences of AVUT compared with examples of existing benchmarks are shown in Table 1.
AVUT introduces a suite of carefully designed audio-centric video understanding tasks, which cover not only content understanding of the video but also the interaction between audio and visual information, such as time synchronization or content matching. Besides, AVUT adopts a dual-dataset design which combines high-quality human annotation with scalable, semi-automatic annotation to obtain a holistic evaluation benchmark. Specifically, a novel Gemini-powered in-context learning data creation approach is proposed to generate high-quality question-answer pairs that are further validated by human annotators. AVUT incorporates 2,662 videos meticulously selected from YouTube, covering 18 audio-centric domains. This collection is split into two partitions: (1) AV-Human, comprising 698 videos with 1,734 expertly human-annotated question-answer pairs; and (2) AV-Gemini, containing the remaining 1,964 videos with 9,875 question-answer pairs obtained from the proposed in-context learning data creation method. In addition, to address the text shortcut problem, a novel answer permutation-based filtering mechanism is proposed and applied in AVUT, ensuring the necessity of multimodal inputs. The main contributions are as follows:
  • We introduce AVUT, the Audio-centric Video Understanding benchmark without Text shortcut, featuring diverse tasks encompassing both audio content understanding and audio-visual alignment.
  • AVUT first systematically investigate the text shortcut problem in existing audio and video benchmarks, and proposes a novel filtering mechanism to minimize the text shortcut and ensure sufficient necessity of the multimodal inputs.
  • We evaluate a diverse range of audio-visual, visual-only, and audio-only models on AVUT to demonstrate their capabilities on audio-centric video understanding.
@@ 2 Related Work
Early research in multimodal alignment explored the creation of shared representations across images, audio, and text Aytar et al. (2017), demonstrating the potential of transferable classifiers. Further research on shared representations across modalities leads to encoders that can align with LLMs like Whisper Radford et al. (2023) or CLIP Radford et al. (2021), making MLLMs technically possible. Recent progress in audio-visual learning has spurred the development of large language models (LLMs) capable of processing both audio and visual information. Early video understanding works like VideoChat Li et al. (2023a) uses external subtitles for speech. Subsequent models have explored approaches to fuse audio and visual modalities. PandaGPT Su et al. (2023) integrates ImageBind Girdhar et al. (2023) and Vicuna Chiang et al. (2023), leveraging image-text pairings for enhanced cross-modal understanding, while VAST Chen et al. (2023b), trained on the large-scale VAST-27M Chen et al. (2023b) dataset, further advances multimodal capabilities. Models like CAT Ye et al. (2025) employ specialized architectures, such as clue aggregators, to effectively process audio-visual events. Video-LLaMA Zhang et al. (2023), VideoLLaMA2 Cheng et al. (2024) and Baichuan Omni Li et al. (2024b) have been fine-tuned specifically on audio data to improve multimodal comprehension. Architectures like the multi-resolution causal Q-former in video-SALMONN Sun et al. (2024) aim to bridge pre-trained audio-visual encoders with LLMs efficiently. Furthermore, models like NExTGPT Wu et al. (2024), OneLLM Han et al. (2024), and VITA Fu et al. (2024b) use “omni-encoders” to uniformly process various data modalities, including audio and video.
@@
Numerous efforts have focused on developing video benchmarks Xiao et al. (2021); Li et al. (2021, 2024a); Ning et al. (2023); Li et al. (2023b); Liu et al. (2024); Mangalam et al. (2023); Fu et al. (2024a); Fang et al. (2024), predominantly targeting the visual aspects of video content. However, the crucial role of audio in video understanding has received comparatively less attention. Several early datasets explore the use of audio in video understanding. AVSD Alamri et al. (2019) focuses on scene-aware dialogue, Pano-AVQA Yun et al. (2021) on panoramic videos, MUSIC-AVQA Li et al. (2022) specifically on musical contexts, and VGGSound Chen et al. (2020) on short videos of audio events. Thus, these datasets target specific scenarios. While AVQA Yang et al. (2022) explores the synergy between audio and visual information but is limited by shorter video lengths and simplistic task design. More recent datasets like UnAV-100 Geng et al. (2023), VALOR Chen et al. (2023a), and VAST-27M Chen et al. (2023b) offer large-scale data but prioritize specific tasks, such as audio event detection, retrieval, and captioning. Similarly, LongVALE Geng et al. (2024) focuses on audio-visual-language events, and AVHBench Sung-Bin et al. (2024) on cross-modal hallucination. These benchmarks, often lacking human annotation, are less suited for evaluating the nuanced and complex aspects of audio-centric understanding, particularly alignment capabilities, targeted by AVUT.
@@ 3.1 Data Curation Principles
Focusing on Audio-Centric Understanding in Video. Unlike existing video benchmarks that predominantly target visual understanding, AVUT focuses specifically on the crucial role of audio in video comprehension. To this end, the tasks in AVUT are designed to focus on the following two aspects: (1) Audio Content Understanding, requiring models to perceive and understand audio information; (2) Audio Visual Alignment, requiring the models to align audio with the corresponding visual information.
Covering Diverse Audio-Centric Video Domains and Comprehensive Annotation. To thoroughly evaluate audio-centric understanding capabilities in MLLMs, AVUT incorporates a diverse selection of 18 video domains where audio plays a central role. For robust evaluation, AVUT offers both AV-Human, a meticulously human-annotated dataset, and AV-Gemini, a significantly larger dataset generated via a novel Gemini-augmented approach. This combination of diverse content and rich annotation enables AVUT to effectively support the research in audio-centric video understanding.
@@ 3.2 Audio-Centric Task Definition
Table 2: Task examples of AVUT. “Audio Visual Segment Matching” is to sort four randomly shuffled audio segments in a video into their original order. “Audio Visual Speaker Diarization” is to transcribe the speech of a visually specified speaker with the Word Error Rate (WER) as the metric.
@@ 3.3 Dataset Construction
AVUT contains 2,662 videos with 11,609 questions in total with each question either created or validated by human annotators. This section describes the data construction in detail.
@@ 3.3.1 Video Collection
We define 18 audio-centric video domains (see Appendix A) to cover a diverse range of scenarios in which audio plays a crucial role in comprehension. Our annotators searched YouTube for videos within these audio-centric domains to manually collect a total of 2,662 raw videos, as depicted in Figure 1. The collected videos were required to be in English, rich in audio content, and recently published (mostly after 2023) to minimize overlap with existing training sets. Videos were intentionally limited to two minutes as driven by the following considerations: First, the nature of our audio-centric tasks, which focus on analyzing specific audio events or aligning audio cues with corresponding visual information within localized segments, lends itself well to evaluation within this shorter time frame. Second, most popular open-source MLLMs struggle to process longer videos, and hence short videos allow us to perform a wider range of comparisons. Finally, extending the length of the video would lead to an exponential increase in the cost and potential inaccuracies of manual annotation, without significantly enhancing the discriminatory power of our benchmark. These videos further went through a validation process, where independent annotators confirmed the designated domains and the presence of sufficient and relevant audio information. The detailed video collection process is provided in Appendix D.
@@ 3.3.2 Human Annotation Partition
Figure 1: The Pipeline of AVUT. The pipeline consists of three main parts: video collection, human annotation and in-context learning creation. Annotators select videos from YouTube and these videos are then used for human annotation, where question-answer pairs are designed based on various task types. The remaining videos are used for in-context learning with Gemini 1.5 Pro. The in-context learning process involves four stages: (1) Demonstration Selection; (2) Prompt Designing; (3) Generation; (4) Refining.
@@
Annotation Procedure. We selected 698 videos for full annotation where annotators were given detailed task definitions and annotation examples. Then, for each video, annotators selected the three most suitable task types and created corresponding multiple-choice questions and answer options, resulting in a total of 1,734 question-answer pairs. For AVDiar, annotators selected 260 videos from the full annotation set that contain a higher number of speakers (3.59 speakers per video on average). Detailed annotations for these videos included speaker appearance, start and end times of utterances, and speech transcriptions. The Audio-Visual Matching task did not require manual annotation, as audio-video segment pairings were generated automatically using a script.
Quality Assurance. For each task in AVUT, we selected and trained annotators with relevant expertise. A rigorous two-stage quality control process, involving specialist review and cross-validation, ensured high-quality annotations. During annotation, questions were carefully designed to require genuine multimodal understanding, avoiding common-sense reasoning and text-based answers.
@@ 3.3.3 Gemini-Augmented Partition
Stage 3: Automated Annotation Generation. We provided Gemini with the remaining 1,964 raw videos from our collected set of 2,662. For each video and each task type, Gemini generates a question, four options and the correct option, resulting in a total of 9,875 generated QA pairs.
Stage 4: Human Review and Refinement. All generated QA pairs underwent meticulous human review, focusing on accuracy, relevance, adherence to task requirements, and answer correctness. Annotators manually corrected any discrepancies. (See Appendix F for correction examples.)
4 Text Shortcut Problem
Text shortcut occurs when models give correct answers by only looking at the textual content of the question without requiring the corresponding video, which results in unreliable evaluation of the video understanding capabilities. This section analyzes the causes and consequences of text shortcut and details the methodology employed in AVUT to mitigate this issue.
Text shortcut arises from several factors. Bias in question formulation can inadvertently provide textual cues or biases that hint at the correct answer, even without viewing the video. For instance, a question about a "fast-moving animal" might bias the model towards choosing "cheetah" even if the video depicts a different animal. Furthermore, models may have encountered similar question-answer pairings during their training, effectively memorizing the associations rather than learning to reason about video content. Finally, the powerful text processing capabilities of advanced language models like GPT-4o allow them to exploit subtle textual cues that humans annotators might overlook.
Text shortcut severely compromises evaluation validity. Models may appear to perform better than they genuinely do, masking their true video understanding capabilities and leading to inflated performance metrics. This can, in turn, misdirect research efforts towards optimizing for textual cues rather than genuine video comprehension, hindering true progress in multimodal learning.
AVUT adopts a filtering mechanism based on answer permutation. For each multiple-choice question, we created four cyclic permutations of the answer options (ABCD, BCDA, CDAB, DABC). The permutations help exclude the influence of potential positional bias in the LLM. We then present these variations of the same question to GPT-4o and check the correctness under each permutation. Without video inputs, if the model can still find the correct answer under all four permutations, the question is likely to have a text shortcut. These questions, deemed solvable through textual cues alone, were subsequently removed from AVUT. This filtering mechanism ensures that the video content is essential to finding the correct answer in the remaining questions.
@@ 5 Experiment
Models AIE (%) ACC (%) AEL (%) AVCM (%) AVOM (%) AVTM (%) Audio Visual MLLMs Gemini 1.5 Pro (2024)     90.02 / 94.70 50.85 / 39.99 83.43 / 73.32 74.05 / 83.13 74.79 / 76.62 81.22 / 83.68 VideoLLaMA2 (7B) (2024)     42.30 / 35.65 32.77 / 32.88 32.02 / 25.75 46.80 / 57.41 49.60 / 46.11 48.92 / 40.90 video-SALMONN (13B) (2024)     52.45 / 32.86 33.33 / 30.17 36.05 / 32.48 34.47 / 41.42 30.25 / 31.42 39.95 / 36.21 PandaGPT (13B) (2023)     24.94 / 30.82 25.00 / 29.13 26.16 / 30.61 24.31 / 27.54 26.28 / 28.44 25.77 / 26.36 Visual MLLMs GPT-4o (2024)     60.37 / 61.43 35.00 / 31.14 33.14 / 34.57 51.81 / 55.71 59.34 / 55.43 70.69 / 61.71 Qwen2-VL (7B) (2024)     59.21 / 53.48 41.67 / 33.18 34.88 / 39.44 62.90 / 65.08 55.85 / 57.01 69.74 / 58.37 LLaVA-Video (7B) (2024)     54.19 / 55.77 39.05 / 30.62 24.83 / 38.17 62.83 / 62.53 61.87 / 58.21 62.94 / 60.04 InternVL2 (76B) (2024)     53.15 / 45.49 25.83 / 26.83 30.23 / 26.89 58.42 / 57.79 53.18 / 50.13 61.70 / 53.25 InternVL2 (8B) (2024)     44.52 / 37.70 30.00 / 25.24 29.07 / 26.58 49.25 / 53.54 47.43 / 43.38 53.19 / 43.69 VILA-1.5 (8B) (2024)     41.72 / 38.28 28.33 / 26.78 28.49 / 36.61 53.30 / 62.38 46.41 / 47.77 46.34 / 46.64 VideoLLaVA (7B) (2023)     34.73 / 27.97 15.00 / 19.10 22.67 / 17.30 36.67 / 42.36 34.09 / 35.86 35.93 / 29.69 Audio MLLM SALMONN (13B) (2024)     45.92 / 45.18 33.33 / 31.68 29.65 / 31.56 32.62 / 32.70 36.76 / 33.11 34.52 / 38.39
@@
Models AV-Human AV-Gemini Overall Audio Visual MLLMs Gemini 1.5 Pro (2024)     78.34     75.21     75.67 VideoLLaMA2 (7B) (2024)     44.90     39.75     40.56 video-SALMONN (13B) (2024)     38.33     34.08     34.74 PandaGPT (13B) (2023)     25.38     28.83     28.31 Visual MLLMs GPT-4o (2024)     56.62     50.00111We evaluated GPT-4 on a subset of 2100 questions randomly sampled from the full question set with balanced task types, sufficiently representing the original task distribution.     53.31 Qwen2-VL (7B) (2024)     58.38     51.05     52.26 LLaVA-Video (7B) (2024)     56.52     50.62     51.48 InternVL2 (76B) (2024)     52.62     43.29     44.72 InternVL2 (8B) (2024)     45.90     38.31     39.47 VILA-1.5 (8B) (2024)     44.48     43.05     43.27 VideoLLaVA (7B) (2023)     33.14     28.70     29.37 Audio MLLM SALMONN (13B) (2024)     36.48     35.43     35.59
Table 4: Overall Performance on AVUT. This table presents the overall performance of each model on the AV-Human, AV-Gemini, and Overall (combined AV-Human and AV-Gemini) datasets. Performance is measured by accuracy (%) on MCQs.